evaluate (-6c)^2 and 6c^2 when c = 10
Answer by stanbon(75887)

evaluate (-6c)^2 and 6c^2 when c = 10
-----
(-6c)^2 = (-6*10)^2 = (-60)^2 = 3600
-------------------------------
6c^2 = 6*(10)^2 = 6*100 = 600
================================
Cheers,
Stan H.